OpenStudy (moonlight93):

I have the answers, but I need to know how to get it. I need help. In a sample distribution, x=56 corresponds to z= 1.00 and x=47 corresponds to z = -0.50. find the mean and standard deviation for the sample. ____________ Answer: mean= 50 standard deviation= 6

OpenStudy (marco26):

I wonder if you still need the solution, but here it is: \[z=\frac{ (x-mean) }{ SD }\] where SD is the standard deviation. when x=56 and z=1: \[1=\frac{ 56-mean }{ SD }\] \[SD= 56- mean \] --> eq. 1 when x=47, z=-0.5: \[-0.5=\frac{ 47-mean }{ SD }\] but we have an equation for SD previously, substitute it \[-0.5=\frac{ 47-mean}{ 56-mean }\] Solving for mean, you will get 50. Substitute 50 to get SD: \[SD=56-mean=56-50=6\]

OpenStudy (moonlight93):

Yes I do, I'm studying for an upcoming exam :) Thank you! @macro26